OpenStudy (anonymous):

1) Simplify into one term        x x x x y x y x y cos 2 sin 2 sin sec cot cos 90 sin 90 sin         

OpenStudy (anonymous):

Please rewrite the question, that isn't readable.

OpenStudy (anonymous):

awkward ok sorry :(

OpenStudy (unklerhaukus):

for the first one \[\square ~\cdot~\square=\square^2\]

OpenStudy (unklerhaukus):

\[\square \cdot\square \cdot\square \cdot\square\cdot \square\cdot x\cdot x\cdot x\cdot x\cdot\square \cdot y=\square^{6}\cdot x^4\cdot y\]